# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ 7.7 Finite-dimensional verification of the generalized eigenproblem
Acknowledgements: The research was in part supported by the NSF grants DMS-2205590 and DMS-2512878. We would like to thank Jiajie Chen, Jia Hao, Xuefeng Liu, Xiang Qin, and Vladimir Sverak for a number of stimulating discussions.
@@ Appendix A Sharp Sobolev Constant
By Cauchy-Schwarz, we have
@@ Appendix B Estimates of Finite Rank Approximation
and we thus conclude the proof by Cauchy-Schwarz. ∎
@@
By $Q$-orthogonality, we have the estimate using (B.2), (4.16), and (B.3) that
@@
where $\mu=\frac{K_{6}}{K_{5}(1-\varepsilon_{4})^{2}}$ is chosen to balance the coefficients of the two terms using $E$-orthogonality, leveraging the weighted AM-GM inequality.
For $I_{3}$, we expand it into double integrals and use Cauchy-Schwarz to get
@@ Appendix D Expression of Operators in Spherical Coordinates
For axisymmetric scalar $f$, our velocity vector $\widetilde{U}$, and pressure $\widetilde{P}$, we know we have the following operators in spherical coordinates:
@@
$\widetilde{\partial}_{\beta}f$ $:=$ $\sin\beta(\cos\beta\partial_{\beta}f-\sin\beta f),$
$\widetilde{\nabla}_{\operatorname{sph}}\widehat{U}$ $:=$ $\frac{\cos^{2}\beta}{\sin\beta}\left(\begin{array}[]{ccc}\widetilde{\partial}_{\beta}\widehat{U}_{r}&\widetilde{\partial}_{\beta}\widehat{U}_{\theta}&\widetilde{\partial}_{\beta}\widehat{U}_{\phi}\\ \partial_{\theta}\widehat{U}_{r}-\widehat{U}_{\theta}&\partial_{\theta}\widehat{U}_{\theta}+\widehat{U}_{r}&\partial_{\theta}\widehat{U}_{\phi}\\ -\widehat{U}_{\phi}&-\widehat{U}_{\phi}\cot\theta&\widehat{U}_{\theta}\cot\theta+\widehat{U}_{r}\end{array}\right),$
@@
$\widetilde{\nabla}_{\operatorname{sph}}\widehat{P}$ $:=$ $\frac{\cos^{2}\beta}{\sin\beta}(\widetilde{\partial}_{\beta}\widehat{P},\partial_{\theta}\widehat{P},0),$
$\widetilde{\operatorname{div}}_{\operatorname{sph}}\widehat{U}$ $:=$ $\frac{\cos^{2}\beta}{\sin\beta}\left(\frac{1}{2}\sin 2\beta\partial_{\beta}\widehat{U}_{r}+\frac{1}{2}(3+\cos 2\beta)\widehat{U}_{r}+\frac{1}{\sin\theta}\partial_{\theta}(\widehat{U}_{\theta}\sin\theta)\right).$
Appendix E Spherical Harmonics
For a scalar function $p$, we can expand $p$ with scalar spherical harmonics:
$p(r,\omega)=\sum_{l=0}^{+\infty}\sum_{m=-l}^{l}p_{lm}(r)S_{lm}(\omega),\quad r\geqslant 0,\quad\omega\in\mathbb{S}^{2}.$
where
$Y_{lm}(\omega)=S_{lm}(\omega)\vec{e}_{r},\quad X_{l}(\omega)=\nabla_{\mathbb{S}^{2}}S_{lm}(\omega),\quad Z_{lm}(\omega)=\vec{e}_{r}\times\nabla_{\mathbb{S}^{2}}S_{lm}(\omega).$
Since we only consider axisymmetric functions, we choose $m=0$ and $\omega$ reduce to $\theta$:
$Y_{l}(\theta)=S_{l}(\theta)\vec{e}_{r},\quad X_{l}(\theta)=S_{l}^{\prime}(\theta)\vec{e}_{\theta},\quad Z_{l}(\theta)=\vec{e}_{\phi}S_{l}^{\prime}(\theta),\quad S_{l}(\theta)=P_{l}(\cos\theta),$
We have the reduced expansions
We note that $X_{l},Y_{m},Z_{n}$ are orthogonal with one another for all subscripts $l,m,n$. Furthermore, they are normalized in the $L^{2}$ inner product:
The gradients of $p$ and $u$ have the following forms, by Appendix D and the eigenvector properties of spherical harmonics and Legendre polynomials:
$\nabla_{\operatorname{sph}}p$ $=$ $\sum_{l=0}^{+\infty}\left(p_{l}^{\prime}(r)Y_{l}(\theta)+\frac{1}{r}p_{l}(r)X_{l}(\theta)\right),$ (E.1)
$\Delta_{\operatorname{sph}}p$ $=$ $\sum_{l=0}^{+\infty}\left(\Delta_{r}p_{l}(r)-\frac{\mu_{l}}{r^{2}}p_{l}(r)\right)S_{l}(\theta),$ (E.2)
$\operatorname{div}_{\operatorname{sph}}u$ $=$ $\sum_{l=0}^{+\infty}\left(\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}+\frac{2}{r}u_{l}^{Y}-\frac{\mu_{l}}{r}u_{l}^{X}\right)S_{l},$ (E.3)
$\Delta_{\operatorname{sph}}u$ $=$ $\sum_{l=0}^{+\infty}(\Delta_{r}u_{l}^{Y}Y_{l}+\Delta_{r}u_{l}^{X}X_{l}+\Delta_{r}u_{l}^{Z}Z_{l})$ (E.4)
where we define
Appendix F Eigenfunction in the Ball
@@
where we integrated by parts in the second step. The corresponding eigenvalue problem reads:
$\Delta_{\operatorname{sph}}u^{b}+\nabla_{\operatorname{sph}}p^{b}=\lambda^{b}u^{b},\quad\operatorname{div}_{\operatorname{sph}}u^{b}=0,$ (F.1)
@@
We expand $u^{b}$ and $p^{b}$ in vector spherical harmonics as follows:
@@
From (F.1), taking the divergence yields $\Delta_{\operatorname{sph}}p=0$. Thanks to (E.2), we obtain for each $l$,
@@
The regular solution is $p_{l}(r)=C_{l}r^{l}$, while the singular branch $r^{-l-1}$ is discarded. Substituting into (E.1)(E.4), the eigenvalue problem becomes
$-\Delta_{r}u_{l}^{Y}-\frac{1}{r^{2}}(2\mu_{l}u_{l}^{X}-(\mu_{l}+2)u_{l}^{Y})+lC_{l}r^{l-1}=\lambda u_{l}^{Y},$ (F.2)
@@
Using (E.3), the divergence-free condition reduces to
$\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}+\frac{2}{r}u_{l}^{Y}-\frac{\mu_{l}}{r}u_{l}^{X}=0.$ (F.3)
@@
We first solve for $u_{l}^{Y}$ and $u_{l}^{X}$. When $l=0$, the divergence-free condition (F.3) forces $u^{Y}_{l}=0$. Thus, we only consider $l\geqslant 1$. Plugging (F.3) into (F.2), we obtain:
$\frac{\mathrm{d}^{2}}{\mathrm{d}r^{2}}u_{l}^{Y}-\frac{4}{r}\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}+\frac{1}{r^{2}}(l(l+1)-2)u_{l}^{Y}+lC_{l}r^{l-1}=\lambda u_{l}^{Y}.$ (F.4)
@@
It is not hard to check that the pair indeed solves (F.2).
@@
$\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}(1)-p_{l}(1)=B_{l}k\tilde{j}_{l}^{\prime}(k)-(k^{2}-l(l-1))A_{l}=0.$ (F.5)
@@
where we use that $\tilde{j}_{l}(kr)$ is a homogeneous solution to (F.4). By (F.5), this becomes
@@
$(k^{2}-l(l-1)(l+2))k\tilde{j}_{l}^{\prime}(k)+(k^{2}-l(l-1))(k^{2}-(l-1)(l+2))\tilde{j}_{l}(k)=0.$ (F.6)
@@
$\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Z}(1)=kJ_{l+\frac{1}{2}}^{\prime}(k)-\frac{1}{2}J_{l+\frac{1}{2}}(k)=0,$ (F.7)
@@
and $R_{l}(\tilde{\beta}_{l,j})$ is finite. Therefore, for every $\tilde{\beta}_{l,j}$, in both cases we obtain we obtain
@@
At $\tilde{\beta}_{l,0}=0$, Taylor expansion using the power series expansion
$J_{l+\frac{1}{2}}(x)=\frac{(x/2)^{l+\frac{1}{2}}}{\Gamma(l+\frac{3}{2})}(1-\frac{x^{2}}{4(l+\frac{3}{2})}+O(x^{4})),$
@@
Therefore, $F_{l}(k)>0$ for sufficiently small $k$. We conclude that $F_{l}(k)$ has at least one root on every interval $(\tilde{\beta}_{l,j},\tilde{\beta}_{l,j+1})$ with $j\geqslant 0$. The case $l=1$ follows by a similar argument.
@@
Appendix G Proof of Propositions 3 and 4
@@
$\tfrac{1}{2}U_{\mathrm{near}}-\tfrac{1}{2}\,\xi\cdot\nabla U_{\mathrm{near}}+\Pi(\widetilde{U}\cdot\nabla\widetilde{U})-\Delta U_{\mathrm{near}}=F,\quad\nabla\cdot U_{\mathrm{near}}=0,$ (G.1)
@@
$F:=\tfrac{1}{2}U_{\mathrm{far}}+\tfrac{1}{2}\,\xi\cdot\nabla U_{\mathrm{far}}+\Delta U_{\mathrm{far}},\quad\nabla^{k}F(\xi)=O(|\xi|^{-3-k})\ \ (k\geqslant 0).$ (G.2)
@@
$u^{\epsilon}=\rho_{\epsilon}*U_{\operatorname{near}},\quad F^{\epsilon}=\rho_{\epsilon}*F.$ (G.3)
@@
$\frac{1}{2}u^{\epsilon}-\frac{1}{2}\xi\cdot\nabla u^{\epsilon}-\frac{1}{2}[\rho_{\epsilon}*,\xi\cdot\nabla]U_{\operatorname{near}}+\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U})-\Delta u^{\epsilon}=F^{\epsilon},$ (G.4)
@@
$\|\nabla^{k-1}[\rho_{\epsilon}*,\xi\cdot\nabla]U_{\operatorname{near}}\|_{L^{2}}\leqslant\|\rho_{\epsilon}(y)y\|_{L^{1}}\|\nabla^{k}U_{\operatorname{near}}\|_{L^{2}}\lesssim\epsilon\|U_{\mathrm{near}}\|_{H^{k}}<\|U_{\mathrm{near}}\|_{H^{k}},\quad\forall k>0.$ (G.5)
We recall two basic facts that will be used repeatedly: the boundedness of Leray projection in $L^{2}$, and that mollification is bounded in $H^{k}$ via Young’s inequality: for any integer $k\geqslant 0$,
$\|\nabla^{k}u^{\epsilon}\|_{L^{2}}=\|\rho_{\epsilon}*\nabla^{k}U_{\mathrm{near}}\|_{L^{2}}\leqslant\|\rho_{\epsilon}\|_{L^{1}}\|\nabla^{k}U_{\mathrm{near}}\|_{L^{2}}=\|\nabla^{k}U_{\mathrm{near}}\|_{L^{2}}.$ (G.6)
Next, we will obtain an estimate of the $H^{k-1}$-seminorm of the nonlinear term. Recall $U_{\operatorname{far}}\in W^{\infty,\infty}$ and $\nabla U_{\operatorname{far}}\in H^{\infty}$. A canonical term in the expansion of $)\nabla^{k-1}(\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U}$)) has the form $\rho_{\epsilon}*\Pi(\nabla^{i}U_{\operatorname{s}}\otimes\nabla^{k-i}U_{\operatorname{t}})$, where $\operatorname{s},\operatorname{t}\in\{\operatorname{far},\operatorname{near}\}$, $0\leqslant i<k$. By Young’s inequality, all of the terms where at least one of $\operatorname{s},\operatorname{t}$ is $\operatorname{far}$ have $L^{2}$-norm bounded by $1+\|U_{\operatorname{near}}\|_{H^{k}}$. When $\operatorname{s},\operatorname{t}=\operatorname{near}$ and $i>0$, we have by Young’s inequality, Holder’s inequality, and Sobolev embedding that
$\|\rho_{\epsilon}*\Pi(\nabla^{i}U_{\operatorname{s}}\otimes\nabla^{k-i}U_{\operatorname{t}})\|_{L^{2}}\leqslant\|\nabla^{i}U_{\operatorname{near}}\|_{L^{4}}\|\nabla^{k-i}U_{\operatorname{near}}\|_{L^{4}}\lesssim\|U_{\operatorname{near}}\|_{H^{k}}^{2}.$ (G.7)
@@
where we use the Cauchy-Schwarz and (G.6) in the first inequality, the variance identity $E_{y}(y-E_{y})^{2}=1/2E_{y,z}(y-z)^{2}$ in the second equality, and the difference-quotient bound $\|u(\cdot+h)-u(\cdot)\|\leqslant|h|\|\nabla u\|$ (via the Lebiniz’s rule and Jensen’s inequality) in the second inequality.
Combined with (G.7), we already collect a good nonlinear estimate for $k>1$ by Sobolev embedding and Young’s inequality:
$\|\nabla^{k-1}\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U})\|_{L^{2}}\lesssim 1+\|U_{\operatorname{near}}\|_{H^{k}}^{2}.$ (G.8)
When $k=1$, we need to treat the term $\|u^{\epsilon}\|_{L^{\infty}}$ with special care using Sobolev embedding as
@@
where we use the interpolation inequality and (G.6) again. Hence, we conclude
$\|\nabla^{k-1}\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U})\|_{L^{2}}\lesssim 1+\|U_{\operatorname{near}}\|_{H^{k}}^{2}+\|\nabla^{k+1}u^{\epsilon}\|_{L^{2}}^{3/4}\|U_{\operatorname{near}}\|_{H^{k}}^{5/4},\quad k=1.$ (G.9)
We are now ready to perform an induction on $k$ to show that $U_{\operatorname{near}}\in H^{k}$. It holds for $k=0,1$, and suppose it holds for any index no greater than $k$, at level $k>0$ we take derivatives of (G.4):
$\frac{1+k}{2}\nabla^{k}u^{\epsilon}-\frac{1}{2}(\xi\cdot\nabla)\nabla^{k}u^{\epsilon}-\frac{1}{2}\nabla^{k}([\rho_{\epsilon}*,\xi\cdot\nabla]U_{\operatorname{near}})+\nabla^{k}(\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U}))-\Delta\nabla^{k}u^{\epsilon}=\nabla^{k}F^{\epsilon},$ (G.10)
@@
In the same fashion, the commutator term enjoys the estimate by (G.5)
@@
and the nonlinear term enjoys a similar estimate by (G.8) and (G.9)
@@
Again, by Young’s inequality, the right hand side of (G.10) has the trivial estimate
@@
Step 1: $L^{2}$-level estimate, that $|\xi|U_{\operatorname{near}},|\xi|\nabla U_{\operatorname{near}}\in L^{2}$. Recall that $U_{\operatorname{near}}$ is smooth, and (G.1) holds in a strong sense with $F\in H^{\infty}$, with decay at the infinity $\nabla^{k}F=O(|\xi|^{-3-k})$. We rewrite (G.1) in the velocity-pressure formation
$\frac{1}{2}U_{\operatorname{near}}-\frac{1}{2}\xi\cdot\nabla U_{\operatorname{near}}+\widetilde{U}\cdot\nabla\widetilde{U}+\nabla P-\Delta U_{\operatorname{near}}=F,$ (G.11)
@@
$\Delta P=-\operatorname{tr}((\nabla\widetilde{U})^{2}),\quad P=\sum_{i,j}R_{i}R_{j}(\widetilde{U}_{i}\widetilde{U}_{j}).$ (G.12)
@@
It holds that $\phi_{L}=|\xi|^{2}$ if $|\xi|<L$, $\phi_{L}=|2L|^{2}$ if $|\xi|>2L$, $\phi_{L}$ is radially increasing since $\chi_{L}$ is radially decreasing, and that $|\Delta\phi_{L}|\lesssim 1$ is uniformly bounded in $L$ due to the estimates on derivatives of $\chi_{L}$. Now we take inner products of (G.11) with $\chi_{R}\phi_{L}U_{\operatorname{near}}$, where the cutoff radius $R$ is much larger than $L$, and we get the following estimates as we take $R\to\infty$.
@@
The pressure term is controlled by (G.12) using Calderón–Zygmund theory $\|P\|\lesssim\|\widetilde{U}\|^{2}_{L^{4}}\lesssim 1$:
@@
The forcing term has the trivial estimate by Cauchy-Schwarz

### v2

@@ 7.5 $L^{\infty}$ bounds for Bessel functions
Acknowledgements: The research was in part supported by the NSF grants DMS-2205590 and DMS-2512878. We would like to thank Jiajie Chen, Joel Dahne, Jia Hao, Xuefeng Liu, Xiang Qin, and Vladimir Sverak for a number of stimulating discussions.
@@ Appendix A Sharp Sobolev Constant
By the Cauchy-Schwarz inequality, we have
@@ Appendix B Estimates of Finite Rank Approximation
and we thus conclude the proof by the Cauchy-Schwarz inequality. ∎
@@
By the $Q$-orthogonality, we have the estimate using (B.2), (4.16), and (B.3) that
@@
where $\mu=\frac{K_{6}}{K_{5}(1-\varepsilon_{4})^{2}}$ is chosen to balance the coefficients of the two terms using the $E$-orthogonality, leveraging the weighted AM-GM inequality.
For $I_{3}$, we expand it into double integrals and use the Cauchy-Schwarz inequality to get
@@ Appendix D Expression of Operators in Spherical Coordinates
For axisymmetric scalar $f$, our velocity vector $\widetilde{U}$, and pressure $\widetilde{P}$, we know we have the following operators in the spherical coordinates:
@@
$\widehat{\partial}_{\beta}f$ $:=$ $\sin\beta(\cos\beta\partial_{\beta}f-\sin\beta f),$
$\widehat{\nabla}_{\operatorname{sph}}\widehat{U}$ $:=$ $\frac{\cos^{2}\beta}{\sin\beta}\left(\begin{array}[]{ccc}\widehat{\partial}_{\beta}\widehat{U}_{r}&\widehat{\partial}_{\beta}\widehat{U}_{\theta}&\widehat{\partial}_{\beta}\widehat{U}_{\phi}\\ \partial_{\theta}\widehat{U}_{r}-\widehat{U}_{\theta}&\partial_{\theta}\widehat{U}_{\theta}+\widehat{U}_{r}&\partial_{\theta}\widehat{U}_{\phi}\\ -\widehat{U}_{\phi}&-\widehat{U}_{\phi}\cot\theta&\widehat{U}_{\theta}\cot\theta+\widehat{U}_{r}\end{array}\right),$
@@
$\widehat{\nabla}_{\operatorname{sph}}\widehat{P}$ $:=$ $\frac{\cos^{2}\beta}{\sin\beta}(\widehat{\partial}_{\beta}\widehat{P},\partial_{\theta}\widehat{P},0),$
$\widehat{\operatorname{div}}_{\operatorname{sph}}\widehat{U}$ $:=$ $\frac{\cos^{2}\beta}{\sin\beta}\left(\frac{1}{2}\sin 2\beta\partial_{\beta}\widehat{U}_{r}+\frac{1}{2}(3+\cos 2\beta)\widehat{U}_{r}+\frac{1}{\sin\theta}\partial_{\theta}(\widehat{U}_{\theta}\sin\theta)\right).$
Appendix E Eigenfunction in the Ball
@@
where we integrate by parts in the second step. The corresponding eigenvalue problem reads:
$\Delta_{\operatorname{sph}}u^{b}+\nabla_{\operatorname{sph}}p^{b}=\lambda^{b}u^{b},\quad\operatorname{div}_{\operatorname{sph}}u^{b}=0,$ (E.1)
@@
We expand $u^{b}$ and $p^{b}$ in the vector spherical harmonics as follows:
@@
From (E.1), taking the divergence yields $\Delta_{\operatorname{sph}}p=0$. Thanks to (5.12), we obtain for each $l$,
@@
The regular solution is $p_{l}(r)=C_{l}r^{l}$, while the singular branch $r^{-l-1}$ is discarded. Substituting into (5.11)(5.14), the eigenvalue problem becomes
$-\Delta_{r}u_{l}^{Y}-\frac{1}{r^{2}}(2\mu_{l}u_{l}^{X}-(\mu_{l}+2)u_{l}^{Y})+lC_{l}r^{l-1}=\lambda u_{l}^{Y},$ (E.2)
@@
Using (5.13), the divergence-free condition reduces to
$\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}+\frac{2}{r}u_{l}^{Y}-\frac{\mu_{l}}{r}u_{l}^{X}=0.$ (E.3)
@@
We first solve for $u_{l}^{Y}$ and $u_{l}^{X}$. When $l=0$, the divergence-free condition (E.3) forces $u^{Y}_{l}=0$. Thus, we only consider $l\geqslant 1$. Plugging (E.3) into (E.2), we obtain:
$\frac{\mathrm{d}^{2}}{\mathrm{d}r^{2}}u_{l}^{Y}-\frac{4}{r}\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}+\frac{1}{r^{2}}(l(l+1)-2)u_{l}^{Y}+lC_{l}r^{l-1}=\lambda u_{l}^{Y}.$ (E.4)
@@
It is not hard to check that the pair indeed solves (E.2).
@@
$\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Y}(1)-p_{l}(1)=B_{l}k\tilde{j}_{l}^{\prime}(k)-(k^{2}-l(l-1))A_{l}=0.$ (E.5)
@@
where we use that $\tilde{j}_{l}(kr)$ is a homogeneous solution to (E.4). By (E.5), this becomes
@@
$(k^{2}-l(l-1)(l+2))k\tilde{j}_{l}^{\prime}(k)+(k^{2}-l(l-1))(k^{2}-(l-1)(l+2))\tilde{j}_{l}(k)=0.$ (E.6)
@@
$\frac{\mathrm{d}}{\mathrm{d}r}u_{l}^{Z}(1)=kJ_{l+\frac{1}{2}}^{\prime}(k)-\frac{1}{2}J_{l+\frac{1}{2}}(k)=0,$ (E.7)
@@
and $R_{l}(\tilde{\beta}_{l,j})$ is finite. Therefore, for every $\tilde{\beta}_{l,j}$, in both cases we obtain
@@
At $\tilde{\beta}_{l,0}=0$, the power series expansion
$J_{l+\frac{1}{2}}(x)=\frac{(x/2)^{l+\frac{1}{2}}}{\Gamma(l+\frac{3}{2})}(1-\frac{x^{2}}{4(l+\frac{3}{2})}+O(x^{4}))$
@@
Therefore, we obtain $F_{l}(k)>0$ for sufficiently small $k$. We conclude that $F_{l}(k)$ has at least one root on every interval $(\tilde{\beta}_{l,j},\tilde{\beta}_{l,j+1})$ with $j\geqslant 0$. The case $l=1$ follows by a similar argument.
@@
Appendix F Proof of Propositions 3 and 4
@@
$\tfrac{1}{2}U_{\mathrm{near}}-\tfrac{1}{2}\,\xi\cdot\nabla U_{\mathrm{near}}+\Pi(\widetilde{U}\cdot\nabla\widetilde{U})-\Delta U_{\mathrm{near}}=F,\quad\nabla\cdot U_{\mathrm{near}}=0,$ (F.1)
@@
$F:=\tfrac{1}{2}U_{\mathrm{far}}+\tfrac{1}{2}\,\xi\cdot\nabla U_{\mathrm{far}}+\Delta U_{\mathrm{far}},\quad\nabla^{k}F(\xi)=O(|\xi|^{-3-k})\ \ (k\geqslant 0).$ (F.2)
@@
$u^{\epsilon}=\rho_{\epsilon}*U_{\operatorname{near}},\quad F^{\epsilon}=\rho_{\epsilon}*F.$ (F.3)
@@
$\frac{1}{2}u^{\epsilon}-\frac{1}{2}\xi\cdot\nabla u^{\epsilon}-\frac{1}{2}[\rho_{\epsilon}*,\xi\cdot\nabla]U_{\operatorname{near}}+\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U})-\Delta u^{\epsilon}=F^{\epsilon},$ (F.4)
@@
$\|\nabla^{k-1}[\rho_{\epsilon}*,\xi\cdot\nabla]U_{\operatorname{near}}\|_{L^{2}}\leqslant\|\rho_{\epsilon}(y)y\|_{L^{1}}\|\nabla^{k}U_{\operatorname{near}}\|_{L^{2}}\lesssim\epsilon\|U_{\mathrm{near}}\|_{H^{k}}<\|U_{\mathrm{near}}\|_{H^{k}},\quad\forall k>0.$ (F.5)
We recall two basic facts that will be used repeatedly: the boundedness of the Leray projection in $L^{2}$, and that mollification is bounded in $H^{k}$ via Young’s inequality: for any integer $k\geqslant 0$,
$\|\nabla^{k}u^{\epsilon}\|_{L^{2}}=\|\rho_{\epsilon}*\nabla^{k}U_{\mathrm{near}}\|_{L^{2}}\leqslant\|\rho_{\epsilon}\|_{L^{1}}\|\nabla^{k}U_{\mathrm{near}}\|_{L^{2}}=\|\nabla^{k}U_{\mathrm{near}}\|_{L^{2}}.$ (F.6)
Next, we will obtain an estimate of the $H^{k-1}$-seminorm of the nonlinear term. Recall $U_{\operatorname{far}}\in W^{\infty,\infty}$ and $\nabla U_{\operatorname{far}}\in H^{\infty}$. A canonical term in the expansion of $)\nabla^{k-1}(\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U}$)) has the form $\rho_{\epsilon}*\Pi(\nabla^{i}U_{\operatorname{s}}\otimes\nabla^{k-i}U_{\operatorname{t}})$, where $\operatorname{s},\operatorname{t}\in\{\operatorname{far},\operatorname{near}\}$, $0\leqslant i<k$. By Young’s inequality, all of the terms where at least one of $\operatorname{s},\operatorname{t}$ is $\operatorname{far}$ have $L^{2}$-norm bounded by $1+\|U_{\operatorname{near}}\|_{H^{k}}$. When $\operatorname{s},\operatorname{t}=\operatorname{near}$ and $i>0$, we have by Young’s inequality, Holder’s inequality, and the Sobolev embedding that
$\|\rho_{\epsilon}*\Pi(\nabla^{i}U_{\operatorname{s}}\otimes\nabla^{k-i}U_{\operatorname{t}})\|_{L^{2}}\leqslant\|\nabla^{i}U_{\operatorname{near}}\|_{L^{4}}\|\nabla^{k-i}U_{\operatorname{near}}\|_{L^{4}}\lesssim\|U_{\operatorname{near}}\|_{H^{k}}^{2}.$ (F.7)
@@
where we use the Cauchy-Schwarz inquality and (F.6) in the first inequality, the variance identity $E_{y}(y-E_{y})^{2}=1/2E_{y,z}(y-z)^{2}$ in the second equality, and the difference-quotient bound $\|u(\cdot+h)-u(\cdot)\|\leqslant|h|\|\nabla u\|$ (via the Lebiniz’s rule and Jensen’s inequality) in the second inequality.
Combined with (F.7), we already collect a good nonlinear estimate for $k>1$ by the Sobolev embedding and Young’s inequality:
$\|\nabla^{k-1}\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U})\|_{L^{2}}\lesssim 1+\|U_{\operatorname{near}}\|_{H^{k}}^{2}.$ (F.8)
When $k=1$, we need to treat the term $\|u^{\epsilon}\|_{L^{\infty}}$ with special care using the Sobolev embedding as
@@
where we use the interpolation inequality and (F.6) again. Hence, we conclude
$\|\nabla^{k-1}\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U})\|_{L^{2}}\lesssim 1+\|U_{\operatorname{near}}\|_{H^{k}}^{2}+\|\nabla^{k+1}u^{\epsilon}\|_{L^{2}}^{3/4}\|U_{\operatorname{near}}\|_{H^{k}}^{5/4},\quad k=1.$ (F.9)
We are now ready to perform an induction on $k$ to show that $U_{\operatorname{near}}\in H^{k}$. It holds for $k=0,1$, and suppose it holds for any index no greater than $k$, at level $k>0$ we take derivatives of (F.4):
$\frac{1+k}{2}\nabla^{k}u^{\epsilon}-\frac{1}{2}(\xi\cdot\nabla)\nabla^{k}u^{\epsilon}-\frac{1}{2}\nabla^{k}([\rho_{\epsilon}*,\xi\cdot\nabla]U_{\operatorname{near}})+\nabla^{k}(\rho_{\epsilon}*\Pi(\widetilde{U}\cdot\nabla\widetilde{U}))-\Delta\nabla^{k}u^{\epsilon}=\nabla^{k}F^{\epsilon}.$ (F.10)
@@
In the same fashion, the commutator term enjoys the estimate by (F.5)
@@
and the nonlinear term enjoys a similar estimate by (F.8) and (F.9)
@@
Again, by Young’s inequality, the right hand side of (F.10) has the trivial estimate
@@
Step 1: $L^{2}$-level estimate, that $|\xi|U_{\operatorname{near}},|\xi|\nabla U_{\operatorname{near}}\in L^{2}$. Recall that $U_{\operatorname{near}}$ is smooth, and (F.1) holds in a strong sense with $F\in H^{\infty}$, with decay at the infinity $\nabla^{k}F=O(|\xi|^{-3-k})$. We rewrite (F.1) in the velocity-pressure formation
$\frac{1}{2}U_{\operatorname{near}}-\frac{1}{2}\xi\cdot\nabla U_{\operatorname{near}}+\widetilde{U}\cdot\nabla\widetilde{U}+\nabla P-\Delta U_{\operatorname{near}}=F,$ (F.11)
@@
$\Delta P=-\operatorname{tr}((\nabla\widetilde{U})^{2}),\quad P=\sum_{i,j}R_{i}R_{j}(\widetilde{U}_{i}\widetilde{U}_{j}).$ (F.12)
@@
It holds that $\phi_{L}=|\xi|^{2}$ if $|\xi|<L$, $\phi_{L}=|2L|^{2}$ if $|\xi|>2L$, $\phi_{L}$ is radially increasing since $\chi_{L}$ is radially decreasing, and that $|\Delta\phi_{L}|\lesssim 1$ is uniformly bounded in $L$ due to the estimates on derivatives of $\chi_{L}$. Now we take inner products of (F.11) with $\chi_{R}\phi_{L}U_{\operatorname{near}}$, where the cutoff radius $R$ is much larger than $L$, and we get the following estimates as we take $R\to\infty$.
@@
The pressure term is controlled by (F.12) using Calderón–Zygmund theory $\|P\|\lesssim\|\widetilde{U}\|^{2}_{L^{4}}\lesssim 1$:
@@
The forcing term has the trivial estimate by the Cauchy–Schwarz inequality